# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v7

@@ Abstract
Tokenization is a necessary component within the current architecture of many language models, including the transformer-based large language models (LLMs) of Generative AI, yet its impact on the model’s cognition is often overlooked. We argue that LLMs demonstrate that the Distributional Hypothesis (DH) is sufficient for reasonably human-like language performance (particularly with respect to inferential lexical competence), and that the emergence of human-meaningful linguistic units among tokens and current structural constraints motivate changes to existing, linguistically-agnostic tokenization techniques, particularly with respect to their roles as (1) vehicles for conveying salient distributional patterns from human language to the model and as (2) semantic primitives. We explore tokenizations from a BPE tokenizer; extant model vocabularies obtained from Hugging Face and tiktoken; and the information in exemplar token vectors as they move through the layers of a RoBERTa (large) model. Besides creating sub-optimal semantic building blocks and obscuring the model’s access to the necessary distributional patterns, we describe how tokens and pretraining can act as a backdoor for bias and other unwanted content, which current alignment practices may not remediate. Additionally, we relay evidence that the tokenization algorithm’s objective function impacts the LLM’s cognition, despite being arguably meaningfully insulated from the main system intelligence, and discuss implications for architectural choices and our understanding of meaning construction.
@@ 2.1 Exemplar tokenizations
As a familiar starting place, we divided the classic Alice in Wonderland novel Carroll (1865) into rough sentence chunks using the simplest possible (imperfect) strategy (splitting on punctuation). The resulting example first chunk has the words “bank” and “the” in it, which we return to later in Sec. 2.3 (this excerpt shown in Fig. 2). For greater linguistic generality, we also used the larger wikitext dataset Merity et al. (2016) for training the tokenizer; in that case, for consistency, we added the extracted first chunk from Alice in Wonderland. [^30] So that, in both cases, we could be certain that the algorithm had the example sentence in the corpus it used to create the vocabulary.
@@ 2.2 Exemplar vocabularies
Many of the tokens that were common amongst the initial batch of vocabularies were also in GPT-4o’s vocabulary, which is worth noting as it suggests that the de facto tokenization approach has not been entirely upended in the time since we started this project (see Fig. 11).
@@
Along the same lines, we noticed that many tokens are cased (or otherwise minor) variations of each other. There were also quite a few tokens allocated to formatting (see Figure 6). Other potentially unhelpful [^42] From the perspective of an English-language LLM; of course, there are contexts in which at least some of these tokens could be useful, such as in a multilingual or non-English LLM! tokens can be seen in Fig. 32: some such tokens represent non-English words or punctuation characters and appeared in exactly one source file. Numbers were not uncommon tokens either, but because the tokenization strategy treats all symbols alike, the number tokens do not directly reflect our conventions (e.g., the base 10 digits, maybe additional milestones like powers of 10, etc.); rather, they reflect common sequences of numbers, which could come from e.g., telephone numbers, backend software development, website URLs, etc.
@@ 2.3 Extispicy: exemplar tokens
We passed contiguous spans of text (a context window of 100 tokens on either side of the token of interest) sourced from a relatively small corpus of English-language fiction through a RoBERTa (large) Masked Language Model (MLM); we parsed the source text into tokens using a BPE tokenizer, the same approach RoBERTa uses. We then extracted the latent space data for our token of interest (in this paper, “ bank”, Fig. 9, “ run”, Fig. 40, and “ the”, Fig. 41, with leading spaces) by saving the embedding vector output by each layer of the model. We used consolidation strategies—UMAP and PCA—to boil down the large amount of data into something human-readable (through truncating and clustering). UMAP can result in “finer clustering than is necessarily present in the data,” which should be kept in mind when interpreting the resultant clusters McInnes, Healy, and Melville (2020, ). For more details of this process, see Sec. 25, Sec. 42. Our reasons for using the RoBERTa model are discussed in the Appendix (Sec. 7). These particular gnogeographic maps are unique to RoBERTa and the text we chose to use; other models and other contexts would both yield different maps. What we are exploring is what these internal organizations can look like, and what kind of strategies a model can use to organize. Although the model’s architecture (including its training objective) and how it was trained will each impact the resultant gnogeographic maps, given that LLMs as a group exhibit impressive natural language competency compared to older models and do have significant commonalities between them, we don’t see any reason to think these gnogeographic maps would be irrelevant outside of this specific combination of model and corpus.
@@
Polysemous words were immediately visually distinguishable from more monosemous words. [^52] Highly polysemous “set” and “run” have many small, well-separated blobs floating around along with a few larger, more meandering clusters that represent the most common meanings. The majority of the clusters are very small and compact. There are shape structures that seem to be in conversation with each other, mainly visible within or from the larger clusters (Fig. 40). On the other hand, “teacher”, a word often considered fairly monosemous, has what look like three very large main clusters, two of which look to be connected in content by thin lines of utterances. The shape of all three could be related to each other. The few small clusters outside of the main three generally look like they may be in conversation with the main ones (in terms of their shapes and direction). See Fig. 42. We can clearly see “naturally interpretable clusters of closely related concepts” within polysemous words, as has been described among human language in lexical networks Youn et al. .
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^53] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions545454e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Fig. 41).
There looked to be both inter- and intra- cluster structure, with identifiable sections of specificity (as to both word meaning and syntax) within a cluster with a broader theme, and across or between clusters. There can be fine-grained distinctions such as between relational, descriptive, restrictive, and attributive roles, as well as distinctions in construction, and possibly even between idiomatic vs. productive uses (Fig. 9).
@@
Some of the clusters seem to reflect putatively fundamental human-like distinctions such as prototypical usage, embodiment, and concreteness (relevant to the Swadesh list and processes of metaphorical extension Maudslay et al. (2024); Swadesh (1952)). [^57] On the left in Fig. 9, a small column of clusters suggests organization around embodiment–biological components and action defined relative to a body (“[food, blood, gene] bank” and “banked” as in turn, respectively). Around the lower center, there seems to be a larger area involving significant figurative components, e.g., “piggybank” and idiomatic uses. We also saw a cluster in “run” focused on its temporal ending (as in stop, e.g., “run into a wall”; Aktionsart), which is notable since time and space are core dimensions to the human experience and universally made salient by the technology of language Rijkhoff (2002). Speculatively, this could be caused by a kind of indirect grounding in human experience (perhaps not unlike what can happen for people when they receive diegetic approximations for what they are not directly experiencing Liu, van Paridon, and Lupyan (2025)). [^58] The potential ‘ceiling’ for what can be conveyed in language depends (at least) on both the properties of language and the architecture of the speakers. This is being borne out in emerging research with deep learning models, exploring under what circumstances their representations diverge or converge with human representations (as can be captured by limited metrics) Xu et al. (2025); Zhuang, Fedorenko, and Andreas (2024). That LLMs could be sensitive to concepts of time, space, and even bodies, is particularly notable since the embodiment [^59] Bodies also being the primary mediator through which we perceive time and space. of an LLM is strikingly inhuman. These clusters are also suggestive of a porous boundary between world knowledge and linguistic knowledge, since they reflect a human world model and human experience through something like word meanings, which could be both promising for LLM gnogeography and informative as to how language is implemented for people (as well as hinting at potential alternatives).
@@ 3.1 Meaning and grounding
3.2 Objective functions
As one of the main forces shaping their cognition, [^76] For an example of the deep-rooted impact of objective function on cognition, Wolfe and Caliskan, 2022, found that “CLIP embeddings show that the high anisotropy observed by Ethayarajh (2019) is not the inevitable result of contextualization, nor even of a specific language modeling architecture, but is connected to the pretraining objective of the model” Wolfe and Caliskan (2022). (They mean pretraining as in before fine-tuning, so they still mean the objective function of the main intelligence.) The importance of the objective function to model cognition is also evident in results from transfer learning. exploring the space of objective functions is promising for allowing LLMs to get more out of the same data [^77] As in Zhuang et al., 2024, who propose that greater model efficiency could be achieved through more “cognitively aligned language learning technologies”, and found improvements in certain language representations when they modified the objective functions to include visual grounding Zhuang, Fedorenko, and Andreas (2024). and to solve some of the shortcomings that seem to come from such a stringent emphasis on producing plausible language above all else, a strategy that makes current LLMs silvertongued but ultimately unreliable and unhumanlike in terms of any behaviour outside those bounds (e.g., this could include aspects of Grice’s maxims, moral consistency, reciprocal self-disclosure, etc.) Grice (1989); Nunes et al. (2024).
@@ 3.2 Objective functions
We show that the tokenization objective function can impact the internal representations of LLMs, not only in choice of tokens (which is self-evident) but in the content of the tokens as well. By forcing orthographically related terms (e.g. ‘run’/‘rung’, Fig. 40) to share a significant portion of their initial embedding information, token-based co-lexification induces representational challenges that the model must handle, meaning that tokenization’s objective function – here, BPE’s orthographic-efficiency-maximizing objective – can shape the model’s internal representations. Of note, the pathway of influence here is arguably at greater remove from the core intelligence than the aforementioned optimizer or the model’s objective function.
In Fig. 40, there are clusters for ‘rung’ and ‘runnels’ encoded via the token ‘ run’. The semantic information from instances using the words ‘runnel’ and ‘rung’ are only co-tokenified with run in the model because of the tokenization objective function of orthographic efficiency agnostic of meaning. That ‘run’ is a component of the strings ‘runnel’ and ‘rung’ is more closely related to their respective orthographies than to their meanings, whether syntagmatic, syntactic, or semantic. The model endeavors to place token vectors meaningfully with respect to other token vectors within its latent space, so the inclusion of a new sense or meaning impacts the placement of other vectors.
Thus the choice of receptacle, of atomic unit, for linguistic representation matters for eventual semantic representations: an emergent connection between Harris’ “purely linguistic” meaning and the koinos kosmos (human world model). (Recall Section 3.1.) [^78] The word ‘runnel’ could be tokenized any number of ways when encountered in input. For example, run+[n+el,ne+l,n+e+l,nel] or r+un, etc. As aforementioned, the tokens available in the vocabulary can obscure the salience of the necessary distributions (Fig. 2, Fig. 3), but the number of possible ways an utterance could be tokenized when the model encounters it also contributes to this obfuscation (tokenization is deterministic but context-dependent). Both aspects of tokenization, the choice of vocabulary pieces and the chunking of input and output, are implicated. The more possible tokenizations given the vocabulary, the more patterns have to be reconciled by the model. Considered from an MDL perspective, this is suspect.
@@ 3.3 Flattening
The tokens form the semantic primitives for the model. They likely need to be well-formed receptacles for the kind of semantic information that makes up human world models, as that is a primary role of words for us.797979“A primary function of the lexicon is to encode novel concepts and ideas as they emerge over time” Xu et al. (2024). The Sapir-Whorf hypothesis in general, nameability (Lupyan’s Label-feedback hypothesis), and even theories from language acquisition like fast-mapping offer explanations for what kind of processes might need such structures Scholz et al. (2024); Lupyan (2012); Gelman and Brandone (2010). [^80] Although LLMs have something like MVP language, how that maps to the intersubjective or intrasubjective competence humans derive from language is unknown.
@@
3.4 Two strategies
Huang et al., 2023’s results Huang et al. (2023b) hint that, not only are symbols and relations the bottom-most necessary components of language, but much of the MVP meaning could be encoded morphosyntactically, and the MVP lexical component could be comparatively minimal, maybe even non-existent. [^86] Maybe the lexical role has to supply any extradiegetic reference, and acts as a diegetic compression mechanism., [^87] However, their results are not straightforward to interpret. Although the tokens are untethered from an initial meaningful embedding with each new context, that doesn’t mean the lexinvariant model didn’t create some representations reasonably equivalent to a lexical component at a higher level, perhaps as features, which it then maps to new symbols depending on the distributions in the new context, especially with the character-level vocabulary. In that case, since a context length of 512 is so few, we think the model could have representations that aid it in mapping between the patterns in the new context and both the distributional patterns it observed and the semantic concepts it built up during training. Otherwise, there would seem to be an extreme poverty-of-the-stimulus problem, a puzzling mismatch between the upper bound of the information in the context and what would be required to explain the decent performance of the model. On the other hand, when the tokens are subword-level, the inner workings of the model could be significantly different, as we think even more of the typical linguistic signal has been disrupted Huang et al. (2023b). This is tentatively reminiscent of theories in which the fundamental unit of meaning construction is relation or metaphor, potentially e.g., theories of C.S. Peirce Peirce (1931), as well as of Hebbian learning. We conjecture that the two meaning-encoding strategies recursively scaffold each other: the compression and referential [^88] Not necessarily as in reference to another mode; this can be reference to another idea or diegetic approximation as well, e.g., jargon. aspect of storing information lexically allow more to be operated on by the relational strategy, [^89] This lexical role would be invaluable from the perspective of information locality. and the compositionality and flexibility of the relational mechanisms allow new variations of the lexical concepts, which can then themselves be named and operated on Xu et al. (2024); Lupyan (2012). Either strategy works for any particular linguistic meaning, but the two together are incredibly powerful—seemingly powerful enough to change gnogeography itself, either by expanding it, or by changing its topography such that distant concepts come to be within arms’ reach.

### v8

@@ Abstract
Tokenization is a necessary component within the current architecture of many language mod-els, including the transformer-based large language models (LLMs) of Generative AI, yet its impact on the model’s cognition is often overlooked. We argue that LLMs demonstrate that the Distributional Hypothesis (DH) is sufficient for reasonably human-like language performance (particularly with respect to inferential lexical competence), and that the emergence of human-meaningful linguistic units among tokens and current structural constraints motivate changes to existing, linguistically-agnostic tokenization techniques, particularly with respect to their roles as (1) vehicles for conveying salient distributional patterns from human language to the model and as (2) semantic primitives. We explore tokenizations from a BPE tokenizer; extant model vocabularies obtained from Hugging Face and tiktoken; and the information in exemplar token vectors as they move through the layers of a RoBERTa (large) model. Besides creating suboptimal semantic building blocks and obscuring the model’s access to the necessary distributional patterns, we describe how tokens and pretraining can act as a backdoor for bias and other unwanted content, which current alignment practices may not remediate. Additionally, we relay evidence that the tokenization algorithm’s objective function impacts the LLM’s cognition, despite being arguably meaningfully insulated from the main system intelligence. Finally, we discuss implications for architectural choices, meaning construction, the primacy of language for thought, and LLM cognition.
@@ 2.1 Exemplar tokenizations
As a familiar starting place, we divided the classic Alice in Wonderland novel Carroll (1865) into rough sentence chunks using the simplest possible (imperfect) strategy (splitting on punctuation). The resulting example first chunk has the words “bank” and “the” in it, which we return to later in Sec. 2.3 (this excerpt shown in Figure 2). For greater linguistic generality, we also used the larger wikitext dataset Merity et al. (2016) for training the tokenizer; in that case, for consistency, we added the extracted first chunk from Alice in Wonderland. [^30] So that, in both cases, we could be certain that the algorithm had the example sentence in the corpus it used to create the vocabulary.
@@ 2.2 Exemplar vocabularies
Many of the tokens that were common amongst the initial batch of vocabularies were also in GPT-4o’s vocabulary, which is worth noting as it suggests that the de facto tokenization approach has not been entirely upended in the time since we started this project (see Figure 11).
@@
Along the same lines, we noticed that many tokens are cased (or otherwise minor) variations of each other. There were also quite a few tokens allocated to formatting (see Figure 6). Other potentially unhelpful [^42] From the perspective of an English-language LLM; of course, there are contexts in which at least some of these tokens could be useful, such as in a multilingual or non-English LLM! tokens can be seen in Figure 32: some such tokens represent non-English words or punctuation characters and appeared in exactly one source file. Numbers were not uncommon tokens either, but because the tokenization strategy treats all symbols alike, the number tokens do not directly reflect our conventions (e.g., the base 10 digits, maybe additional milestones like powers of 10, etc.); rather, they reflect common sequences of numbers, which could come from e.g., telephone numbers, backend software development, website URLs, etc.
@@ 2.3 Extispicy: exemplar tokens
We passed contiguous spans of text (a context window of 100 tokens on either side of the token of interest) sourced from a relatively small corpus of English-language fiction through a RoBERTa (large) Masked Language Model (MLM); we parsed the source text into tokens using a BPE tokenizer, the same approach RoBERTa uses. We then extracted the latent space data for our token of interest (in this paper, “ bank”, Figure 9, “ run”, Figure 40, and “ the”, Figure 41, with leading spaces) by saving the embedding vector output by each layer of the model. We used consolidation strategies—UMAP and PCA—to boil down the large amount of data into something human-readable (through truncating and clustering). UMAP can result in “finer clustering than is necessarily present in the data,” which should be kept in mind when interpreting the resultant clusters McInnes, Healy, and Melville (2020, ). For more details of this process, see Sec. 25, Sec. 42. Our reasons for using the RoBERTa model are discussed in the Appendix (Sec. 6). These particular gnogeographic maps are unique to RoBERTa and the text we chose to use; other models and other contexts would both yield different maps. What we are exploring is what these internal organizations can look like, and what kind of strategies a model can use to organize. Although the model’s architecture (including its training objective) and how it was trained will each impact the resultant gnogeographic maps, given that LLMs as a group exhibit impressive natural language competency compared to older models and do have significant commonalities between them, we don’t see any reason to think these gnogeographic maps would be irrelevant outside of this specific combination of model and corpus.
@@
Polysemous words were immediately visually distinguishable from more monosemous words. [^52] Highly polysemous “set” and “run” have many small, well-separated blobs floating around along with a few larger, more meandering clusters that represent the most common meanings. The majority of the clusters are very small and compact. There are shape structures that seem to be in conversation with each other, mainly visible within or from the larger clusters (Figure 40). On the other hand, “teacher”, a word often considered fairly monosemous, has what look like three very large main clusters, two of which look to be connected in content by thin lines of utterances. The shape of all three could be related to each other. The few small clusters outside of the main three generally look like they may be in conversation with the main ones (in terms of their shapes and direction). See Figure 42. We can clearly see “naturally interpretable clusters of closely related concepts” within polysemous words, as has been described among human language in lexical networks Youn et al. .
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^53] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions545454e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Figure 41).
There looked to be both inter- and intra- cluster structure, with identifiable sections of specificity (as to both word meaning and syntax) within a cluster with a broader theme, and across or between clusters. There can be fine-grained distinctions such as between relational, descriptive, restrictive, and attributive roles, as well as distinctions in construction, and possibly even between idiomatic vs. productive uses (Figure 9).
@@
Some of the clusters seem to reflect putatively fundamental human-like distinctions such as prototypical usage, embodiment, and concreteness (relevant to the Swadesh list and processes of metaphorical extension Maudslay et al. (2024); Swadesh (1952)). [^57] On the left in Figure 9, a small column of clusters suggests organization around embodiment–biological components and action defined relative to a body (“[food, blood, gene] bank” and “banked” as in turn, respectively). Around the lower center, there seems to be a larger area involving significant figurative components, e.g., “piggybank” and idiomatic uses. We also saw a cluster in “run” focused on its temporal ending (as in stop, e.g., “run into a wall”; Aktionsart), which is notable since time and space are core dimensions to the human experience and universally made salient by the technology of language Rijkhoff (2002). Speculatively, this could be caused by a kind of indirect grounding in human experience (perhaps not unlike what can happen for people when they receive diegetic approximations for what they are not directly experiencing Liu, van Paridon, and Lupyan (2025)). [^58] The potential ‘ceiling’ for what can be conveyed in language depends (at least) on both the properties of language and the architecture of the speakers. This is being borne out in emerging research with deep learning models, exploring under what circumstances their representations diverge or converge with human representations (as can be captured by limited metrics) Xu et al. (2025); Zhuang, Fedorenko, and Andreas (2024). That LLMs could be sensitive to concepts of time, space, and even bodies, is particularly notable since the embodiment [^59] Bodies also being the primary mediator through which we perceive time and space. of an LLM is strikingly inhuman. These clusters are also suggestive of a porous boundary between world knowledge and linguistic knowledge, since they reflect a human world model and human experience through something like word meanings, which could be both promising for LLM gnogeography and informative as to how language is implemented for people (as well as hinting at potential alternatives).
@@ 3.1 Meaning and grounding
3.2 Objective functions: POSIWID teleology
As one of the main forces shaping their cognition, [^76] For an example of the deep-rooted impact of objective function on cognition, Wolfe and Caliskan, 2022, found that “CLIP embeddings show that the high anisotropy observed by Ethayarajh (2019) is not the inevitable result of contextualization, nor even of a specific language modeling architecture, but is connected to the pretraining objective of the model” Wolfe and Caliskan (2022). (They mean pretraining as in before fine-tuning, so they still mean the objective function of the main intelligence.) The importance of the objective function to model cognition is also evident in results from transfer learning. exploring the space of objective functions is promising for allowing LLMs to get more out of the same data [^77] As in Zhuang et al., 2024, who propose that greater model efficiency could be achieved through more “cognitively aligned language learning technologies”, and found improvements in certain language representations when they modified the objective functions to include visual grounding Zhuang, Fedorenko, and Andreas (2024). and to solve some of the shortcomings that seem to come from such a stringent emphasis on producing plausible language above all else, a strategy that makes current LLMs silvertongued but ultimately unreliable and unhumanlike in terms of any behaviour outside those bounds (e.g., this could include aspects of Grice’s maxims, moral consistency, reciprocal self-disclosure, etc.) Grice (1989); Nunes et al. (2024). [^78] The section header is in reference to Stafford Beer’s famous coinage, POSIWID, or, “the purpose of a system is what it does”.
@@ 3.2 Objective functions: POSIWID teleology
We show that the tokenization objective function can impact the internal representations of LLMs, not only in choice of tokens (which is self-evident) but in the content of the tokens as well. By forcing orthographically related terms (e.g. ‘run’/‘rung’, Figure 40) to share a significant portion of their initial embedding information, token-based co-lexification induces representational challenges that the model must handle, meaning that tokenization’s objective function – here, BPE’s orthographic-efficiency-maximizing objective – can shape the model’s internal representations. Of note, the pathway of influence here is arguably at greater remove from the core intelligence than the aforementioned optimizer or the model’s objective function.
In Figure 40, there are clusters for ‘rung’ and ‘runnels’ encoded via the token ‘ run’. The semantic information from instances using the words ‘runnel’ and ‘rung’ are only co-tokenified with run in the model because of the tokenization objective function of orthographic efficiency agnostic of meaning. That ‘run’ is a component of the strings ‘runnel’ and ‘rung’ is more closely related to their respective orthographies than to their meanings, whether syntagmatic, syntactic, or semantic. The model endeavors to place token vectors meaningfully with respect to other token vectors within its latent space, so the inclusion of a new sense or meaning impacts the placement of other vectors.
Thus the choice of receptacle, of atomic unit, for linguistic representation matters for eventual semantic representations: an emergent connection between Harris’ “purely linguistic” meaning and the koinos kosmos (human world model). (Recall Section 3.1.) [^79] The word ‘runnel’ could be tokenized any number of ways when encountered in input. For example, run+[n+el,ne+l,n+e+l,nel] or r+un, etc. As aforementioned, the tokens available in the vocabulary can obscure the salience of the necessary distributions (Figure 2, Figure 3), but the number of possible ways an utterance could be tokenized when the model encounters it also contributes to this obfuscation (tokenization is deterministic but context-dependent). Both aspects of tokenization, the choice of vocabulary pieces and the chunking of input and output, are implicated. The more possible tokenizations given the vocabulary, the more patterns have to be reconciled by the model. Considered from an MDL perspective, this is suspect.
@@ 3.3 Flattening
The tokens form the semantic primitives for the model. They likely need to be well-formed receptacles for the kind of semantic information that makes up human world models, as that is a primary role of words for us.808080“A primary function of the lexicon is to encode novel concepts and ideas as they emerge over time” Xu et al. (2024). The Sapir-Whorf hypothesis in general, nameability (Lupyan’s Label-feedback hypothesis), and even theories from language acquisition like fast-mapping offer explanations for what kind of processes might need such structures Scholz et al. (2024); Lupyan (2012); Gelman and Brandone (2010). [^81] Although LLMs have something like MVP language, how that maps to the intersubjective or intrasubjective competence humans derive from language is unknown.
@@
3.4 Two strategies for meaning construction
Huang et al., 2023’s results Huang et al. (2023b) hint that, not only are symbols and relations the bottom-most necessary components of language, but much of the MVP meaning could be encoded morphosyntactically, and the MVP lexical component could be comparatively minimal, maybe even non-existent. [^87] Although the lexical role still has to supply any extradiegetic reference and act as a diegetic compression mechanism., [^88] However, their results are not straightforward to interpret. Although the tokens are untethered from an initial meaningful embedding with each new context, that doesn’t mean the lexinvariant model didn’t create some representations reasonably equivalent to a lexical component at a higher level, perhaps as features, which it then maps to new symbols depending on the distributions in the new context, especially with the character-level vocabulary. In that case, since a context length of 512 is so few, we think the model could have representations that aid it in mapping between the patterns in the new context and both the distributional patterns it observed and the semantic concepts it built up during training. Otherwise, there would seem to be an extreme poverty-of-the-stimulus problem, a puzzling mismatch between the upper bound of the information in the context and what would be required to explain the decent performance of the model. On the other hand, when the tokens are subword-level, the inner workings of the model could be significantly different, as we think even more of the typical linguistic signal has been disrupted Huang et al. (2023b). We hope to see future work extending their research. This is tentatively reminiscent of theories in which the fundamental unit of meaning construction is relation or metaphor, such as those of C.S. Peirce Peirce (1931), or as in Hebbian learning. We could even interpret the overloading of dedicated neurons in metaphorically interpretable ways Forli and Yartsev (2023) as evidence of metaphor, of relation, as the primary mechanism across many organizational scales, from the biologically to the linguistically morphological (for more discussion, see Zimmerman (2025)). We conjecture that the two meaning-encoding strategies recursively scaffold each other: the compression and referential [^89] Not necessarily as in reference to another mode; this can be reference to another idea or diegetic approximation as well, e.g., jargon. aspect of storing information lexically allow more to be operated on by the relational strategy, [^90] This lexical role would be invaluable from the perspective of information locality. and the compositionality and flexibility of the relational mechanisms allow new variations of the lexical concepts, which can then themselves be named and operated on Xu et al. (2024); Lupyan (2012). Either strategy works for any particular linguistic meaning, but the two together are incredibly powerful—seemingly powerful enough to change gnogeography itself, either by expanding it, or by changing its topography such that distant concepts come to be within arms’ reach.